Find differential coefficient dy/dvx if y=epowerx /logx

Grade:12

1 Answers

jitender
6 years ago
y=e^x /logxdy/dx = d[(e^x)/logx]/dx= ~~~> d u/v / x = [vdu/dx - udv/dx]/v^2So dy/dx = [logx (e^x) - e^x (1/x)]/log^2x= e^x [ logx - 1/x]/logx..............And de^x/dx=e^xAnd dlogx/x=1/x